# The Power of Training: How Different Neural Network Setups Influence the Energy Demand

### 03 Jan 2024 v1

@@ Introduction
To compare the results, we generated equation 2 to calculate the breakeven point on how many times the pre-trained model needs to be recycled to compensate its footprint compared to training the full architecture every time.
@@
(a) PAMAP2 on CUDA
(b) PAMAP2 on MPS
Figure 4: Break-even points, highlighting the amount of cycles until the pre-taining energy investment is compensated.
The results are shown in figure 4 for CUDA and MPS training. In this case, the results are solely based on the frozen encoder, so the optimization process does not consider the encoder anymore. The consumption for the pretraining run is shown in solid lines whereas the basic training without encoder pretraining is a dashed line. The crossing of two lines of the same color indicates the break-even point, which means with how many complete downstream-task training cycles, the investment in the pretraining starts to yield better energy savings in the longer perspective. The ideal situation is that the break-even point appears as early as possible. From the 16 setups utilized, we extracted the best, worst, and average scenarios. As we can see, the blue lines for CUDA and MPS both show a scenario where the pretraining already outperforms the standard training after one recycling step. On the other hand, for the worst case, visualized with orange, it takes up to 43 cycles for the CUDA run and up to 23 cycles on MPS until the spent pretraining energy is compensated. As we can derive from equation 2, the compensation depends on the increased energy consumption of the encoder pretraining and the ratio between training the classifier alone or with the encoder together. On average, the energy is compensated after around 10 to 13 recycling steps of the pre-trained encoder.
Additionally, we investigated the average energy consumption between freezing and unfreezing the encoder after its pretraining run. (Marcelino 2018) Since the optimizer does not update the weights during the backpropagation to finetune the encoder, there is less energy necessary to train the classifier. As shown in figure 3, there is on average an energy saving of 61.5 % for the CUDA experiment and 46.1% on MPS across all 16 setup combinations. Since proper pretraining ensures the required encoders’ behavior, the freezing of weights can be a helpful measure to reduce energy consumption.
@@
Encoder CUDA MPS
Unfreeze 4.26 1.42
Freeze 1.64 0.76
Savings 61.5% 46.1%
Multitask Learning Paradigm Results
(a) CelebA on CUDA
(b) CelebA on MPS
Figure 5: Energy Consumption for training attribute and identity classifier separately and training it through the multitask model.
For the multitask scenario trained on CelebA, we utilize the ResNet18 architecture pre-trained on ImageNet as the encoder. Due to ImageNets size and required training time, we decided to recycle an already existing, pre-trained model. As a side note, the best option to save energy in machine learning is the usage of already existing, pre-trained models to compensate their energy consumption further.
@@
The experiment was conducted through the same procedure as the pretraining scenario, training the three models (attribute, identity, and both together) with the 16 different setups. The results are shown in figure 5. For all tested setups, the green bar for training both classifiers in one go is considerably lower than the cumulative energy for training the attribute and identity separately. Even more, the multitask training results in a lower energy consumption than training one single task classifier.
In order to interpret the results, we further calculated the average energy per epoch across the three scenarios as shown in figure 4. Since the average energy per epoch is almost identical across training attributes, identity, or both, no matter if trained on the CUDA or MPS system, we further calculated the average length of the training in number of epochs. As we can see, for this experiment the number of epochs rules the overall energy consumption whereas the efficiency from energy per epoch is negligible. Since the loss functions are connected during the training in the multitask scenario, the classifiers of attribute and identity are able to cross-regulate each other, which results in a quicker convergence. The result manifests our previous findings for minimizing either one or at best both, the energy per epoch and the overall number of epochs used till convergence.
@@
Type CUDA MPS
Attribute 22.47 / 368 40.87 / 359
Identity 22.46 / 150 41.09 / 161
Both 22.44 / 142 41.36 /145
Table 4: Average Energy per Epoch (mWh) / Average number of Epochs
Conclusion
In conclusion, we investigated the energy consumption of two application scenarios on two different hardware configurations. Across all experiments, we followed the approach of training the selected architectures with variations in batch size and learning rates. For the 16 selected setups, we first evaluated their influence on the efficiency through energy per epoch and further the training duration through the overall energy consumption until the early stopping monitor criteria were met. As a result, the batch size influences the energy efficiency whereas the learning rate determines the training duration. The optimal configuration has to maximize the efficiency per epoch and minimize the training duration by setting the optimal batch size and learning rate. If this is the case, such optimizations can lead to energy savings of up to 80% in our tests.
For the learning paradigm testings, the calculated break-even point is an indicator of how and at which recycling iteration the initial energy investment for the pretraining is compensated. A long recycling process can be necessary if too much energy is invested into the pretraining or if the savings are too weak for each recycling step. Lastly, for the multitask training, we not only showed that training two classifiers in one training process is less energy-consuming, we additionally showed that fewer epochs are necessary due to the mutual training support of the classifiers.

### 15 Mar 2024 v2

@@ 5 Pretraining Learning Paradigm Results
To compare the results, we generated equation 2 to calculate the breakeven point on how many times the pre-trained encoder model needs to be recycled in order to compensate the pretraining energy overhead compared to training the full architecture every time.
@@
(a) PAMAP2 on Nvidia H100
(b) PAMAP2 on Nvidia A100
(c) PAMAP2 on Nvidia RTX 3090
Figure 5: Break-even points, highlighting the amount of cycles until the pretraining energy investment is compensated.
The results are shown in Figure 5 for the three desired hardware setups. In this case, the results are solely based on the frozen encoder, so the optimization process does not consider the encoder anymore. The consumption for the pretraining run is shown in solid lines whereas the basic training without encoder pretraining is a dashed line. The crossing of two lines of the same color indicates the break-even point, which means with how many complete downstream-task training cycles, the investment in the pretraining starts to yield better energy savings in the longer perspective. The dashed pre-trained lines commonly start at a higher energy level since the starting point represents the pretraining energy investment, whereas the slope of each line represents the addition of the recycling cycle’s training energy over time. The ideal situation is that the break-even point appears as early as possible. From the 16 setups of hyperparameter configurations utilized, we extracted the best, worst, and average scenarios.
As we can see across the three hardware setups, the blue lines show a scenario where the pretraining already outperforms the standard training after two to three recycling steps. On the other hand, for the worst case, visualized with orange, it takes between 30 to 40 cycles until the spent pretraining energy is compensated. As we can derive from equation 2, the compensation depends on the increased energy consumption of the encoder pretraining and the ratio between training the classifier alone or with the encoder together. On average, the energy is compensated after seven to 13 recycling steps of the pre-trained encoder. Interestingly, in Figure 5c, the worst-case scenario with almost 40 recycling steps still consumes less total energy than the average recycling. The reason for that is the focus on cycles instead of energy in this graphic, leading to the approval that even though you can recycle the energy with fewer cycles, the total consumption has to be taken into account as well in order to maintain the life-cycle efficiency.
Additionally, we investigated the average energy consumption between freezing and unfreezing the encoder’s layer weights after its pretraining run. [15] Since the optimizer does not update the weights during the backpropagation to fine-tune the encoder, there is less energy needed to train the full model. As shown in Table 2, average energy saving between 23% to 45% can be achieved across all 16 setup combinations. Since proper pretraining already ensures the required encoders’ behavior, the freezing of weights can be a helpful action to reduce energy consumption, and further energy investment for fine-tuning can be saved.
@@
Encoder H100 A100 RTX3090
Unfreeze 3.89 Wh 4.24 Wh 4.97 Wh
Freeze 6.01 Wh 5.55 Wh 9.13 Wh
Savings 35.2% 23.6% 45.5%
6 Multitask Learning Paradigm Results
For the multitask scenario trained on CelebA, we utilize the ResNet18 architecture pre-trained on ImageNet as the encoder. Due to ImageNets size and required training time, we decided to recycle an already existing, pre-trained model. As a side note, the best option to save energy in machine learning is the usage of already existing, pre-trained models from external, shared sources to compensate their energy investment further.
@@ 6 Multitask Learning Paradigm Results
The experiment was conducted through the same procedure as the previous pretraining scenario, training the three models (attribute, identity, and both together) with the 16 different setups. We calculated the average energy per epoch across the three scenarios as shown in Table 3. Since the average energy per epoch is almost identical across training attributes, identity, or both, with different operating levels across the three hardware configurations, we further calculated the average length of the training in number of epochs. As we can see, for this experiment the number of epochs rules the overall energy consumption whereas the efficiency from energy per epoch is negligible. Since the loss functions are connected during the training in the multitask scenario, the classifiers of attribute and identity are able to cross-regulate each other, which results in a quicker convergence. Instead of recycling already present knowledge as for the pretraining, this experiment benefits from the active exchange of information through the loss function. The result manifests our previous findings for minimizing either one or at best both, the energy per epoch and the overall number of epochs, to properly decrease the energy consumption for the machine learning training.
Table 3: Average Energy per Epoch (mWh) / Average number of Epochs
@@
Type H100 A100 RTX3090
Attribute 47.93 mWh / 358 46.82 mWh / 380 81.41 mWh / 382
Identity 47.26 mWh / 170 45.79 mWh / 163 81.42 mWh / 159
Both 48.34 mWh / 85 45.48 mWh / 86 81.19 mWh / 84
7 Correlation with High-End Consumer Hardware
We transferred our experiment to high-end consumer hardware to compare our findings independently of the utilized HPC cluster. Therefore we ran our experiment scenarios for the CelebA and the PAMAP2 dataset again on a workstation powered by an Nvidia RTX 6000 Ada with 48GB VRAM. Further, to test the independence of CUDA-based systems, we set up a 16-inch Apple Macbook Pro with M1 Max chip with 32-core GPU and 64GB unified memory as a second device. The Nvidia GPU workstation runs on CUDA 12.3 and the Apple device runs on metal performance shader (MPS) for neural network training acceleration. We implemented a custom tracker based on the powermetrics tool of Mac OS to generate the same energy monitoring and data structure as collecting data through Carbontracker.
For the training regime, we could reproduce the influences of batch size on the energy per epoch efficiency metric and the learning rate influence towards However, the findings were somewhat less distinctive and showed higher variations. We attribute this to the missing isolation of the jobs, which are now running on commercial operating systems. Additionally, for the powermetrics-based tracking, it remains unclear how precise the measurements can be compared to the carbontracker measurements. Even though the local hardware operated on different power levels, running the two learning paradigm tests confirmed our findings that pretraining and multitask learning can be utilized to improve life-cycle-oriented efficiency. For pretraining, the average breakeven points are calculated for 9 cycles on the CUDA setup and 12 cycles on MPS. Finally, for the multitask learning, the information sharing through the loss function by training attribute and identity at the same time could reduce the energy consumption by around 40%.
8 Conclusion
In conclusion, we investigated the energy consumption of two popular ML scenarios on three different HPC and 2 pro-consumer workstation configurations with variations in batch size and learning rates. For the 16 selected setups, we first evaluated their influence on the efficiency through energy per epoch and further, the total energy per training cycle until the early stopping monitor criteria were met. Our findings showcased up to 80% decrease in energy consumption if hyperparameters are set properly. For the learning paradigm tests, the calculated break-even point is an indicator if energy investment for pretraining is beneficial. Proper selection of hyperparameters may help to compensate for the energy overhead already after a few recycling steps. A long recycling process can be necessary if too much energy is invested into the pretraining or if the savings are too weak for each recycling step. Lastly, for the multitask training, we not only showed that training two classifiers in one training process is less energy-consuming, we additionally showed that fewer epochs are necessary in total due to the mutual training support of the classifiers.
As a limitation, we observed that hyperparameters influence the pretraining and downstream task cycles differently, this is mostly determined by the number of epochs needed to reach the accuracy target. Therefore, it is not possible to derive best practice recommendations without trying them out. Moreover, even the dataset influences the energy consumption depending on its size and the chosen number of samples for pretraining and thorough training. Nevertheless, as the goal of training DNNs is to have a model that reaches good performance, optimizing for power consumption alone is not enough. Future research in this direction should balance the model prediction performance and the energy efficiency based on the baseline findings from our work.